OpenStudy (anonymous):

if 12 man can load 4 lorries in 2 hours , find the rate of loading in lorries per men hour?

OpenStudy (anonymous):

\[24\div 6\]

OpenStudy (ash2326):

12 men can load 4 lorries in 2 hours so 1 man can load \(\frac{4}{12}\) lorries in 2 hours in 1 hour one man will load \[\huge{\frac{4}{12}\times \frac{1}{2}=\frac{4}{24}=\frac{1}{6} lorries\ per\ men\ per\ hour}\]